# Uncovering Latent Arguments in Social Media Messaging by Employing LLMs-in-the-Loop Strategy

### v1

@@ 3.5 Mapping Instances to Arguments
Age Group State Entity Talking Points
Table 4: Most mentioned entities and talking points of targeted ads for young, working-age, older population of TX and FL from COVID-19 vaccine campaigns dataset.
4 Case Studies
We explore two case studies involving discussions on social media: (1) climate campaigns and (2) COVID-19 vaccine campaigns. For climate campaigns, we work on the corpus of $14k$ ads released by Islam et al. (2023b). The dataset focuses on climate-related English ads on Facebook based on the US from January $2021$ to January $2022$. All ads in this corpus contain predictions for stance (e.g., pro-energy, clean-energy) and theme (e.g., support climate policy.) (Islam & Goldwasser, 2024). For COVID-19 vaccine campaigns, we use the corpus of $9k$ ads released by Islam & Goldwasser (2022) focusing on COVID-19 vaccine-related English ads on Facebook based on US from December $2020$ to January $2022$. All ads in this corpus contain predictions for moral foundation (e.g., care/harm) (Haidt & Graham, 2007) and theme (e.g., vaccine equity.) (Islam & Goldwasser, 2022). For each ad of both corpora, there are the following attributes: ad ID, title, ad description, ad body, funding entity, spend, impressions, distribution over impressions broken down by gender (male, female, unknown), age (7 groups), and location down to states in the USA. Derived themes are shown in Table 5 in Appendix A.1. Additional details about the dataset can be found in the original publication.
Our main goal is to use our LLMs-in-the-Loop framework to identify prominent arguments in the corpora introduced above. At the beginning, we have ads and their associated themes. Those themes are derived by Islam & Goldwasser (2024; 2022). At the first iteration, we follow the step introduced in Section 3 to obtain the list of talking points under each theme and the mapping $ads\rightarrow arguments$. After iteration $1$, we identify $113$ arguments for Climate campaigns and $47$ arguments for COVID-19 vaccine campaigns (Table 1). In $2^{nd}$ iteration, we repeat this process for the rest of the ads remaining unassigned from iteration $1$ (by applying a threshold of $<0.5$), resulting in the discovery of $100$ and $31$ additional arguments for climate and COVID-19 vaccine campaigns respectively (Table 1). An example of a resulting set of arguments under the ‘Patriotism’ (one of the themes from climate campaigns) and ‘VaccineEquity’ (one of the themes from COVID-19 vaccine campaigns) themes after two rounds of iterations of our LLMs-in-the-Loop are shown in Table 2. The prompt templates are provided in Figure 3 in Appendix A.2. The complete set of final arguments is provided in the Appendix A.3 in Table LABEL:arg_climate and LABEL:arg_covid.
4.1 Evaluation
5 Analysis
5.1 Demographic Targeting
Understanding how messages are tailored to target particular demographic groups is crucial for opinion analysis. To analyze the variation in talking points on climate campaigns across different age groups, we examine targeted ads within three age categories, i.e., young people (ages $13$-$24$), working-age people (ages $25$-$54$), and the older population (age $55+$), focusing on California (CA) and Texas (TX). We choose these states because TX is recognized as the world’s energy capital [^2] https://www.eia.gov/todayinenergy/detail.php?id=49356, while CA has recently positioned itself as a leading advocate in the battle against climate change [^3] https://www.pewtrusts.org/en/research-and-analysis/blogs/stateline/2022/10/06/california-takes-leading-edge-on-climate-laws-others-could-follow. To identify the top-k most frequently mentioned entities in the texts, we prompt LLMs in a zero-shot setting. We show the results in Table 3. We observe that talking points vary by age group in TX. For young people, the focus is on energy reliability and affordability, learning from Texas’ deregulation disaster. For the working-age population, the emphasis is on accelerating the transition to renewable energy to combat climate change and support economic growth through job creation. Meanwhile, for the older population, the discussion centers on the need for urgent legislative action on climate change (Table 3).
(a) Event: Hurricane Ida hits on $29$ August $2021$.
We adopt a similar approach to analyze talking points across three different age groups in COVID-19 vaccine campaigns. We select Florida (FL) because the coronavirus surged to record levels there in August 2021202120212021444https://en.wikipedia.org/wiki/COVID-19_pandemic_in_Florida, and Texas (TX) because, as of April $3^{rd}$, $2021$, its vaccination rates were lagging behind the US average [^5] https://en.wikipedia.org/wiki/COVID-19_pandemic_in_Texas. Results are detailed in Table 4. For the working age group in FL, the ad’s talking point advocates for building and restoring public trust in the COVID-19 vaccine and medical community by criticizing FL Governor Ron DeSantis’s appointment of Dr. Joseph Ladapo as Surgeon General, characterizing Ladapo as an anti-masker and vaccine skeptic. The critique suggests that this appointment is insensitive or disrespectful and emphasizes the need to appoint someone who believes in science, not conspiracy theories. While targeting an older population of FL, the talking point evolves around equitable vaccine distribution and access for senior citizens (Table 4). On the other hand, while targeting a young population of TX, the talking point focuses on COVID-19 vaccine safety for children (Table 4).
5.2 Shifts in Talking Points Following Significant Events
Figure 2(a) displays the top-four shifts in talking points in sponsored advertisements over the five months before and after Hurricane Ida struck the USA. There is a noticeable presence of arguments for ”urgent legislative action on climate change to enhance resiliency against extreme weather events and secure a sustainable future.” as well as advocacy for ”embracing solar energy for its broad benefits, focusing on economic growth, environmental protection, and enhanced accessibility.” following the hurricane (see Figure 2(a)). Figure 2(b) illustrates the talking points used in sponsored ads three days before and after President Biden announced the federal vaccine mandate. In the aftermath, three new predominant arguments emerged: ”Resistance against federal vaccine mandates, emphasizing personal choice and legal action to uphold individual rights.”, ”COVID-19 vaccine mandates, including recent Supreme Court rulings and public opinion on federal mandates.”, ”Against Biden’s vaccine mandate, portraying it as government tyranny and urging resistance and financial support.” (Figure 2(b)). These shifts suggest that sponsored content is highly responsive to real-world events, quickly reflecting and potentially shaping public discourse on pivotal issues.
6 Conclusion
We introduce a simple yet effective iterative LLMs-in-the-Loop framework for uncovering latent arguments from social media messaging. Our quantitative results show that newly discovered talking points can cover a larger portion of messaging and map instances to associated arguments accurately with respect to human judgment. Additionally, our analysis highlights how these talking points are specifically tailored for demographic targeting and dynamically shift in response to real-world events.
7 Ethics Statement
To our knowledge, our research work, as described in this paper, adhered to all ethical guidelines. Instead of introducing a new dataset, we conducted experiments using two existing datasets, which have been properly cited. In this work, we iterated the LLMs-in-the-Loop approach twice. In the future, we aim to determine the optimal number of iterations and identify when to decide to stop exploring unassigned instances. Our approach relies on GPT-4. We use GPT-4 because it requires fewer computational resources compared to open-source alternatives. The author’s personal views are not represented in any qualitative result that we report.
Appendix A Appendix
A.1 Themes
Table 5 shows the derived themes from the climate campaign (Islam & Goldwasser, 2024; Islam et al., 2023b) and COVID-19 vaccine campaign (Islam & Goldwasser, 2022) datasets.
Climate Economy_pro, ClimateSolution, Pragmatism, Patriotism, AgainstClimatePolicy, Economy_clean, FutureGeneration, Environmental, HumanHealth, Animals, SupportClimatePolicy, AltEnergy, PoliticalAffiliation. BidenGasPriceIncrease, AgainstCorporateInterests, GasTax, Deforestation, Carbon, CustomerBasedAltEnergy, EnergyAffordabilityandSustainabilityLegislation, EcofriendlyConsumerChoices, PlasticWasteandEnvironmentalImpact, PromoteSustainableTransportation, WaterManagementandSustainability, FoodSecurity.
Table 5: Derived themes from previous studies: Climate campaigns (Islam & Goldwasser, 2024; Islam et al., 2023b), COVID-19 vaccine campaigns (Islam & Goldwasser, 2022)
A.2 Prompt Templates
In this section, we show the prompt templates used in our work to generate a sub-cluster summary (Figure. 3 (a)), a talking point from each summarized sub-cluster (Figure. 3 (b)), and a refined talking point (Figure. 3 (c)).

### v3

@@ 3.5 Mapping Instances to Arguments
4 Case Studies
We explore two case studies involving discussions on social media: (1) climate campaigns and (2) COVID-19 vaccine campaigns. For climate campaigns, we work on the corpus of $14k$ ads released by Islam et al. (2023b). The dataset focuses on climate-related English ads on Facebook based on the US from January $2021$ to January $2022$. All ads in this corpus contain predictions for stance (e.g., pro-energy, clean-energy) and theme (e.g., support climate policy.) (Islam and Goldwasser, 2024). For COVID-19 vaccine campaigns, we use the corpus of $9k$ ads released by Islam and Goldwasser (2022) focusing on COVID-19 vaccine-related English ads on Facebook based on US from December $2020$ to January $2022$. All ads in this corpus contain predictions for moral foundation (e.g., care/harm) (Haidt and Graham, 2007) and theme (e.g., vaccine equity.) (Islam and Goldwasser, 2022). For each ad of both corpora, there are the following attributes: ad ID, title, ad description, ad body, funding entity, spend, impressions, distribution over impressions broken down by gender (male, female, unknown), age (7 groups), and location down to states in the USA. Derived themes are shown in Table 5 in Appendix A.1. Data statistics regarding the number of instances in each theme for climate (Fig. 4(a)) and COVID-19 vaccine (Fig. 4(b)) campaign are shown in Fig. 4 in App. A.2. Additional details about the dataset can be found in the original publication.
Our main goal is to use our LLMs-in-the-Loop framework to identify prominent arguments in the corpora introduced above. At the beginning, we have ads and their associated themes. Those themes are derived by Islam and Goldwasser (2024, 2022). At the first iteration, we follow the step introduced in Section 3 to obtain the list of talking points under each theme and the mapping $ads\rightarrow arguments$. Hyperparameters are detailed in App. A.3. After iteration $1$, we identify $113$ arguments for Climate campaigns and $47$ arguments for COVID-19 vaccine campaigns (Table 1). In $2^{nd}$ iteration, we repeat this process for the rest of the ads remaining unassigned from iteration $1$ (by applying a threshold of $<0.5$), resulting in the discovery of $100$ and $31$ additional arguments for climate and COVID-19 vaccine campaigns respectively (Table 1). An example of a resulting set of arguments under the ‘Patriotism’ (one of the themes from climate campaigns) and ‘VaccineEquity’ (one of the themes from COVID-19 vaccine campaigns) themes after two rounds of iterations of our LLMs-in-the-Loop are shown in Table 8 in App. A.4. We show the number of sub-clusters identified under each theme for both iterations of the climate (Table 6) and COVID-19 (Table 7) case studies are provided in App. A.2. The prompt templates are provided in Figure 5 in App. A.5. A concrete example of prompt for summarizing top-5 instances under patriotism theme of climate campaign and the generated summary are shown in Fig. 6 in App. A.5. Prompt example of generating talking point from a summary of top-5 instances under patriotism theme of climate campaign dataset is shown in Fig. 7 in App. A.5. The complete set of final arguments is provided in the App. A.6 in Table LABEL:arg_climate and Table LABEL:arg_covid.
4.1 Evaluation
(a) Baseline: 10 LDA Topics.
(b) Baseline: 20 LDA Topics.
(c) Ours: After $2^{nd}$ round of iteration (COVID-19).
(d) Baseline: 10 LDA Topics.
(e) Baseline: 30 LDA Topics.
(f) Ours: After $2^{nd}$ round of iteration (climate).
Figure 2: Correlations between arguments & moral foundations for COVID-19; arguments & stances for climate.
4.2 Argumentative Cohesion Comparison
As there are previously annotated moral foundations for COVID-19 vaccine campaigns (Islam and Goldwasser, 2022) and for climate campaigns, there are annotated energy industry and climate change-related stances (Islam and Goldwasser, 2024; Islam et al., 2023b) available. We use it for baseline analysis. We can characterize arguments/talking points as the reasons people cite to support or oppose contentious topics, such as the climate change debate or the vaccine debate. We consider assignments to be better if they are more cohesive; for example, when arguments are more strongly correlated with specific stances or moral foundations.
For the COVID-19 vaccine debate case study, we calculate the Pearson correlation matrices to evaluate the correlation between arguments and moral foundations, presenting the results in heatmaps (Fig. 2). In Fig. 2(c), we show random $15$ arguments (among $78$ total arguments of COVID-19) to show the correlation between the arguments and the moral foundations. To compare with LDA baseline, we choose $10$ topics (Fig. 2(a)) and $20$ topics (Fig. 2(b)) to keep our $15$ arguments as middle ground. We can interpret reasons as distributions over moral foundations. Our analysis reveals that our arguments (Fig. 2(c)) exhibit higher correlations with moral foundations than those topics derived from LDA (Fig. 2(a), Fig. 2(b)). For example, in Fig. 2(c), we see that liberty/oppression moral foundation is strongly correlated with the talking point “Against COVID vaccine mandates and passports, urging support for personal choice and constitutional rights regarding vaccination.” Additionally, the purity/degradation moral foundation is notably correlated with the talking points “Vaccination as an act of love and protection towards neighbors, aligning with Christian values.” and “Using religion as a means to exempt individuals from mandatory COVID-19 vaccination requirements.” Other expected trends emerge, such as fairness/cheating being highly correlated with talking points related to “Global vaccine equity, emphasizing free and universal access to COVID-19 vaccines to end the pandemic for all”.
For the climate debate case study, we perform a correlation test between the identified talking points and the stance expressed in the ads (i.e., pro-energy or clean-energy). We calculate the Pearson correlation matrices and present them in Fig. 2. We compare our discovered talking points with a set of topics extracted using LDA. In Fig. 2(f), we show random $25$ arguments (among $213$ total arguments of climate) to show the correlation between the arguments and the stance. To compare with LDA baseline, we choose $10$ topics (Fig. 2(d)) and $30$ topics (Fig. 2(e)) to keep our $25$ arguments as middle ground. We observe that our talking points (Fig. 2(f)) have stronger correlations with stances than the derived LDA topics (Fig. 2(d), Fig. 2(e)). For example, we note that pro-energy stance is strongly correlated with talking point such as “Criticizes President Biden’s economic policies for significantly increasing gas prices and living costs, urging a shift in leadership or policy.”
Hyperparameters are detailed in App. A.3.
4.3 Ablation Study
We provide an ablation study in which we mine arguments from the top k instances of a cluster without summarizing them. We find comparable results (Table 2) in terms of coverage based on the distance threshold. We use a threshold determined by the distance from the cluster centroid to an ad, with shorter distances being preferable.
Case Iter. Number of covered ads
Study thr < 0.6 thr < 0.5 thr < 0.4 thr < 0.3
Climate 1 13319 10677 5355 1132
1-w/o sum. 13394 10613 5189 1164
2 13669 11541 6360 1458
2-w/o sum. 13759 11592 6143 1384
COVID-19 1 7962 6525 3589 850
1-w/o sum. 8133 6507 3477 787
2 8197 6833 4015 1089
2-w/o sum. 8426 6767 3710 908
Table 2: Ablation study (coverage). sum: summary, thr: threshold.
Model Acc F1
$Longformer_{text}$ 90.13% 89.89%
$Longformer_{tp}$ 83.43% 83.44%
$\mathbf{Longformer_{text+tp}}$ 93.30% 93.29%
$RoBERTa_{text}$ 93.07% 92.96%
$RoBERTa_{tp}$ 82.73% 82.81%
$\mathbf{RoBERTa_{text+tp}}$ 93.65% 93.56%
$llama3_{text}$ 92.00% 90.95%
$llama3_{tp}$ 81.00% 78.68%
$\mathbf{llama3_{text+tp}}$ 92.50% 91.49%
Table 3: Contribution of talking point (tp) in stance classifier for climate campaigns dataset.
5 Downstream Task
We design a downstream task of stance classification using the uncovered talking points. The task is intuitively designed to learn the association between the stances and the talking points used by the advertisers for contentious debate.
Stance Prediction is designed as a binary classification task. In this task, the model is given a talking point uncovered by our LLMs-in-the-Loop approach and an ad text. The task is to predict the stance of the ad, i.e., pro-energy or clean-energy. We use pre-trained Longformer (Beltagy et al., 2020) model and RoBERTa (Liu et al., 2019) model for the stance classifiers. Besides, we add LLM-based classifier LLaMA (Touvron et al., 2023) for comparison and here we use llama3-70b-8192222https://github.com/meta-llama/llama3. We measure the impact of talking points and present results in Table 3. Our quantitative analysis in Table 3 is designed to demonstrate the contribution of the extracted arguments for the stance identification downstream task. In our ablation study, we compare the performance of classifiers using only text or talking points. We find that the stance prediction task improves when talking points are included with ad text. For example: when using Longformer, macro average F1 score improves from $89.89\%$ (text only) to $93.29\%$; when using RoBERTa, it improves from $92.96\%$ (text only) to $93.56\%$; and when using llama3, it improves from $90.95\%$ (text only) to $91.49\%$ (Table 3). Hyperparameters are provide in App. A.3.
6 Analysis
6.1 Demographic Targeting
Understanding how messages are tailored to target particular demographic groups is crucial for opinion analysis. To analyze the variation in talking points on climate campaigns across different age groups, we examine targeted ads within three age categories, i.e., young people (ages $13$-$24$), working-age people (ages $25$-$54$), and the older population (age $55+$), focusing on California (CA) and Texas (TX). We choose these states because TX is recognized as the world’s energy capital333www.eia.gov//todayinenergy/detail.php?id=49356, while CA has recently positioned itself as a leading advocate in the battle against climate change444www.pewtrusts.org//en/research-and-analysis/blogs/stateline/2022/10/06/california-takes-leading-edge-on-climate-laws-others-could-follow. To identify the top-k most frequently mentioned entities in the texts, we prompt LLMs in a zero-shot setting. We show the results in App. A.8 (Table 9). We observe that talking points vary by age group in TX. For young people, the focus is on energy reliability and affordability, learning from Texas’ deregulation disaster. For the working-age population, the emphasis is on accelerating the transition to renewable energy to combat climate change and support economic growth through job creation. Meanwhile, for the older population, the discussion centers on the need for urgent legislative action on climate change.
(a) Event: Hurricane Ida hits on $29$ August $2021$.
We adopt a similar approach to analyze talking points across three different age groups in COVID-19 vaccine campaigns. We select Florida (FL) because the coronavirus surged to record levels there in August 2021202120212021555https://en.wikipedia.org/wiki/COVID-19_pandemic_in_Florida, and Texas (TX) because, as of April $3^{rd}$, $2021$, its vaccination rates were lagging behind the US average [^6] https://en.wikipedia.org/wiki/COVID-19_pandemic_in_Texas. Results are detailed in Table 4. For the working age group in FL, the ad’s talking point advocates for building and restoring public trust in the COVID-19 vaccine and medical community by criticizing FL Governor Ron DeSantis’s appointment of Dr. Joseph Ladapo as Surgeon General, characterizing Ladapo as an anti-masker and vaccine skeptic. The critique suggests that this appointment is insensitive or disrespectful and emphasizes the need to appoint someone who believes in science, not conspiracy theories. While targeting an older population of FL, the talking point evolves around equitable vaccine distribution and access for senior citizens (Table 4). On the other hand, while targeting a young population of TX, the talking point focuses on COVID-19 vaccine safety for children (Table 4).
Age St. Entity Talking Points
Table 4: Most mentioned entities and talking points of targeted ads for young, working-age, older population of TX and FL from COVID-19 vaccine campaigns dataset. St : State.
6.2 Shifts in Talking Points Following Significant Events
Hurricane Ida: On August $29$, $2021$, Hurricane Ida made landfall near Port Fourchon, Louisiana, as a Category 4 hurricane [^7] https://en.wikipedia.org/wiki/Hurricane_Ida.
Federal Vaccine Mandate: On September $09$, $2021$, President Biden announced an executive order on COVID-19 vaccine mandates for federal workers, large employers, and health care staff888www.whitehouse.gov/briefing-room/presidential-actions/2021/09/09/executive-order-on-requiring-coronavirus-disease-2019-vaccination-for-federal-employees/. Figure 3(a) displays the top-four shifts in talking points in sponsored advertisements over the five months before and after Hurricane Ida struck the USA. There is a noticeable presence of arguments for "urgent legislative action on climate change to enhance resiliency against extreme weather events and secure a sustainable future." as well as advocacy for "embracing solar energy for its broad benefits, focusing on economic growth, environmental protection, and enhanced accessibility." following the hurricane (see Figure 3(a)). Figure 3(b) illustrates the talking points used in sponsored ads three days before and after President Biden announced the federal vaccine mandate. In the aftermath, three new predominant arguments emerged: "Resistance against federal vaccine mandates, emphasizing personal choice and legal action to uphold individual rights.", "COVID-19 vaccine mandates, including recent Supreme Court rulings and public opinion on federal mandates.", "Against Biden’s vaccine mandate, portraying it as government tyranny and urging resistance and financial support." (Figure 3(b)). These shifts suggest that sponsored content is highly responsive to real-world events, quickly reflecting and potentially shaping public discourse on pivotal issues.
7 Conclusion
We introduce a simple yet effective iterative LLMs-in-the-Loop framework for uncovering latent arguments from social media messaging. Our quantitative results show that newly discovered talking points can cover a larger portion of messaging and map instances to associated arguments accurately with respect to human judgment. Besides, our stance prediction downstream task demonstrates that including talking point information helps to improve the performance of the classifier. Additionally, our analysis highlights how these talking points are specifically tailored for demographic targeting and dynamically shift in response to real-world events.
8 Limitations
Though we show the effectiveness of our framework focusing on two case studies, i.e., climate and COVID-19 vaccine campaigns, it is a domain-agnostic framework. However, the main idea of uncovering talking points using LLMs and associating the ad to their closest talking point is applicable to other cases with no changes.
Our approach relies on GPT-4 to generate a multi-document summary and talking point. We chose GPT-4 instead of the open-source counterparts due to computational resource constraints.
LLMs are trained on vast amounts of human-generated text, potentially embedding numerous human biases (Blodgett et al., 2020; Brown et al., 2020). We did not consider any bias in our task.
Some minority viewpoints may not form a cluster or appear in the top-k instances of a cluster to be summarized. In such cases, their viewpoints could be overlooked. In the future, we plan to refine the clusters by integrating techniques like outlier detection (Duan et al., 2009) and feedback mechanisms to better capture and represent minority perspectives in our summaries.
A limitation of this study is the inherent subjectivity in interpreting demographic targeting, as the differentiation of themes based on demographics (specifically age and gender in this paper) is subjective, which can affect the utility of research findings.
9 Ethics Statement
To our knowledge, our research work, as described in this paper, adhered to all ethical guidelines. Instead of introducing a new dataset, we conducted experiments using two existing datasets, which have been properly cited. In this work, we iterated the LLMs-in-the-Loop approach twice. In the future, we aim to determine the optimal number of iterations and identify when to decide to stop exploring unassigned instances. The author’s personal views are not represented in any qualitative result that we report.
Appendix A Appendix
A.1 Themes
Table 5 shows the derived themes from the climate campaign (Islam and Goldwasser, 2024; Islam et al., 2023b) and COVID-19 vaccine campaign (Islam and Goldwasser, 2022) datasets.
Climate Economy_pro, ClimateSolution, Pragmatism, Patriotism, AgainstClimatePolicy, Economy_clean, FutureGeneration, Environmental, HumanHealth, Animals, SupportClimatePolicy, AltEnergy, PoliticalAffiliation, BidenGasPriceIncrease, AgainstCorporateInterests, GasTax, Deforestation, Carbon, CustomerBasedAltEnergy, EnergyAffordabilityandSustainabilityLegislation, EcofriendlyConsumerChoices, PlasticWasteandEnvironmentalImpact, PromoteSustainableTransportation, WaterManagementandSustainability, FoodSecurity.
Table 5: Derived themes from previous studies: Climate campaigns (Islam and Goldwasser, 2024; Islam et al., 2023b), COVID-19 vaccine campaigns (Islam and Goldwasser, 2022)
A.2 Data Statistics
Fig. 4 shows the statistics regarding the number of instances in each theme for climate (Fig. 4(a)) and COVID-19 vaccine (Fig. 4(b)) campaign. In Table 6, we show the number of sub-clusters identified under each theme for both iterations of the climate case. In Table 7, we provide the number of sub-clusters identified under each theme for both iterations of the COVID-19 case study.
(a) Climate campaign.
(b) COVID-19 vaccine campaign.
Figure 4: Number of instances in each theme for climate and COVID-19 vaccine campaign.
Climate theme # sub-clusters per theme (Iter1) # sub-clusters per theme (Iter2)
Economy_pro 6 4
ClimateSolution 3 2
Pragmatism 8 2
Patriotism 4 3
AgainstClimatePolicy 3 2
Economy_clean 4 6
HumanHealth 3 7
FutureGeneration 4 4
Environmental 3 6
Animals 5 5
AltEnergy 5 7
SupportClimatePolicy 7 4
PoliticalAffiliation 5 4
BidenGasPriceIncrease 3 3
AgainstCorporateInterests 7 3
GasTax 7 3
Deforestation 7 6
Carbon 6 3
CustomerBasedAltEnergy 7 5
FoodSecurity 4 5
EnergyAffordabilityandSustainabilityLegislation 4 5
EcofriendlyConsumerChoices 5 3
PlasticWasteandEnvironmentalImpact 6 1
PromoteSustainableTransportation 6 4
WaterManagementandSustainability 4 4
Table 6: Number of sub-clusters per climate theme across two iterations.
COVID-19 theme # sub-clusters per theme (Iter1) # sub-clusters per theme (Iter2)
VaccineRollout 5 2
EncourageVaccination 3 2
VaccineMandate 6 4
VaccineStatus 4 1
GovTrust 2 3
CovidPlan 3 2
VaccineEfficacy 4 2
VaccineMisinformation 2 4
GovDistrust 3 2
VaccineEquity 5 2
VaccineDevelopment 3 N/A
VaccineReligion 2 3
VaccineSymptom 4 4
NaturalImmunity 2 1
Table 7: Number of sub-clusters per COVID-19 theme across two iterations.
@@ A.2 Data Statistics
A.3 Experimental Details
To obtain topics from LDA, we use Gensim Rehurek and Sojka (2011) implementation. We follow the pre-processing steps shown in Hoyle et al. (2021) and estimate the number of topics in a data-driven manner by maximizing. We do a grid search over a set of {10, 20, 25, 30} for the LDA baseline.
For SBERT embedding, we use the sentence transformer model ‘all-mpnet-base-v2’ with default parameters.
For the stance classifier, we randomly split the data into three subsets, namely training set ($60\%$), validation set ($20\%$), and test set ($20\%$). Show the result on the test set in Table 3. For Longformer, we use ‘longformer-base-4096’ model, batch size = $8$, epoch = $3$, learning rate = $52e-$ 5, weight decay = $0.01$. For RoBERTa, we use ‘roberta-base’ model with same batch size, epoch, learning rate, and weight decay used for Longformer. Our early stopping criterion is the lowest validation loss. To run the stance classifier, we use a single GPU GeForce GTX $1080$ Ti GPU, with $4$ Intel Core $i5$-$7400$ CPU @ $3.00$ GHz processors, and it takes around $10$ minutes to run the model.
To run llama3-70b-8192, we use Groq API [^9] https://wow.groq.com/. For tokenizer, we use Meta-Llama-3-70B-Instruct from Hugging Face.
@@ A.3 Experimental Details
A.4 Arguments under specific Themes
Table 8 shows examples of resulting arguments after $1^{st}$ and $2^{nd}$ iteration under the ‘Patriotism’ theme for climate campaigns and the ‘VaccineEquity’ theme for COVID-19 vaccine campaigns theme.